1)prove the identity (sin theta +cos theta)(1 - sin theta cos theta)= sin cube theta + cos
cube theta
2)hence solve the equation (sin theta + cos theta)(1- sin theta cos theta)=3 cos cube theta for 0 degree <=theta<=360 degree
1
Expert's answer
2020-04-21T13:42:08-0400

Given : "(\sin\theta+\cos\theta)(1-\sin\theta\cos\theta)=\sin^3\theta+\cos^3\theta"


Important:

  • Since this is a very basic identity starting on either side is not a big deal but try to identify the easiest side to start.
  • Working out L.H.S to R.H.S needs the application of the identity "\small\cos^2\theta+\sin^2\theta =\small 1" & some simplifications whereas R.H.S to L.H.S needs "a^3+b^3=(a+b)(a^2-ab+b^2)" which gives the answer in seconds.
  • Division by "\cos^3\theta" is possible because the value for theta when "\cos\theta=0" is not a solution for the equation in the 2nd part.



Calculations

1)."\qquad
\begin{aligned}
\small\sin^3\theta+\cos^3\theta&=\small(\sin\theta+\cos\theta)(\sin^2\theta-\sin\theta\cos\theta+\cos^2\theta)\\
&=\small(\sin\theta+\cos\theta)(1-\sin\theta\cos\theta)\cdots\cdots\cdots(\because\sin^2\theta+\cos^2\theta=1)\\
\end{aligned}"

Therefore, the above identity is proved.


2). Now it can be written that, "\small\sin^3\theta+\cos^3\theta = \small 3\cos^3\theta"

"\qquad
\begin{aligned}
\therefore\,\small \sin^3\theta&= \small 2\cos^3\theta\\
\small \tan^3\theta& = \small 2 \cdots\cdots (division \,by\, \cos^3\theta) \\
 \small \tan\theta&= \small 2^{\frac{1}{3}}\\
\small \text{but,} \, \tan\theta&= \small \tan(180+\theta)=\small 2^{\frac{1}{3}}\cdots\cdots(\because\, 0\degree<=\theta<=360\degree)\\
\small \therefore \theta&= \small \tan^{-1}(2^\frac{1}{3})\\
&= \small \bold{51.56\degree}\\
\small 180\degree+\theta &= \small \bold{231.56\degree}\\
\end{aligned}"